HETAV PATEL

Grade 11,

Volume of co2 obtained at STP by the complete decompositions of 9.8g Na2Co3 is

Volume of co2 obtained at STP by the complete decompositions of 9.8g Na2Co3 is

Grade:12th pass

2 Answers

Arun
4 years ago
Na2CO3⇒ Na2O + CO2
1 mole of sodium carbonate give 1 mole of carbon dioxide.
as 105.98g sodium carbonate gives 44.01g carbon dioxide
therefore, 9.85 g sodium carbonate gives 9.85 × 44.01 =4.09 g of CO2 105.98 
Moles of CO2 mole = weight =4.09/44.01= 0.092 mole
Molar mass

1 mole of gas = 22.4 L gas 
Therefore,0.092 mole of CO2 = 22.4 × 0.092 = 2.06 L CO2.
Yashwanth Gowda B D
2 years ago
You may get answer as 2.08 L by calculation. But Na2CO3 is a soda ash which does not decomposes on heating 👍